# Integer Lattice Gas with a sampling collision operator for the fluctuating Navier-Stokes Equation

### v1

@@ IV.1 The Decaying Sound Wave
To test the approximate Boltzmann limit of our lattice gas, we start by analyzing the decaying isothermal sound wave. While the isothermal nature makes this a non-physical application, it is still valuable to compare the lattice Boltzmann and lattice gas algorithms without addressing the complication of including energy conservation. An analytical solution for this system was found by Luiz Czelusniak Czelusniak (2022) in his thesis, and is given by,
@@
$\lambda=\left(\frac{2\tilde{\pi}}{L}\right)^{2}\frac{\nu}{\bar{N}^{eq}},$ (65)
@@
$f^{2}=\left(\frac{2\tilde{\pi}}{L}\right)^{2}\theta^{2}-\lambda^{2},$ (66)
@@
The sound wave must be initialized utilizing a particle distribution for lattice gas systems due to the discrete nature of particles. As the global equilibrium ensemble for the lattice gas is approximately Poisson distributed, we initialize our system by sampling the occupation numbers out of the Poisson distribution from Eq. (31), with $\tilde{f_{i}}$-values given by Eq. (32). The $\bar{N}$- and $u$-values needed to find the $\tilde{f_{i}}$-values can then be defined as the initial conditions for Eqs. (63) and (64) respectively,
@@
Figure 3: Normalized pre- and post-collision $\pi$-values of a decaying sound wave. The lattice Boltzmann system was initialized using Eq. (68) with $\bar{N}^{eq}=1,000$ and $L=250$, and the lattice gas used these same values as a mean for its Poisson distributed initialization. After 50 iterations run at $\omega=1$, the pre- and post-collision $\pi$-values were measured over the course of a single collision. The lattice gas was averaged over 100,000 seeds. Both were normalized according to the equilibrium average post-collision $\pi$-value from Eq. (70).
We can observe the effect of a single collision on the system by looking at the difference between the pre- and post-collision averaged $\pi$-values of the system. We then normalize the system with the expected average post-collision $\pi$-value, which is found by using the $f_{i}^{0}$ from Eq. (32),
@@
To observe the viscosity of the system, we first consider the amplitude decay of the sine wave as it approaches equilibrium. We utilize a method of amplitude extraction discussed by Blommel and Wagner in an appendix to their paper Blommel and Wagner (2018). Specifically, the amplitude of our sound wave is given by,
@@
Figure 4: The extracted amplitude evolution of the decaying sound wave for various values of $\omega$ for the lattice gas and lattice Boltzmann simulations over 850 iterations for a lattice size of 50. Systems were initialized otherwise identically to those in Fig. 3, and run for $\omega=1.0$, $\omega=0.7$, and $\omega=0.1$. The amplitude was extracted utilizing Blommel and Wagner’s method after each iteration Blommel and Wagner (2018). Lattice gas systems were averaged over 25,000 seeds.
Fig. 4 shows the results of this amplitude extraction for $\omega=1$, $\omega=0.7$, and $\omega=0.1$ over 850 iterations for an average of $\bar{N}^{eq}=1,000$ particles. There is excellent agreement between the lattice gas and the lattice Boltzmann algorithms in this case, despite the fact that the BGK lattice Boltzmann algorithm is not an exact Boltzmann limit for our lattice gas. We can measure the viscosity of the system numerically by extracting the decay rate from the wave amplitude. We can do this by fitting our results from Fig. 4 to the analytical solution from Eq. (63). The amplitude of Eq. (63) is given by,
@@
Figure 5: The recovered viscosities of the lattice Boltzmann and lattice Gas corresponding to different values of $\omega$ along with the theoretical scaling from Eq. (61). Each system was initialized and run identically to those in Fig. 4. Along with the systems from Fig. 4, systems were also run at $\omega=0.3$ and $\omega=0.5$. The $\lambda$-values were found via non-linear curve fitting on xmgrace with the analytical solution from Eq. (72) and calculated using Eq. (65).
We can then compare the extracted values of $\nu$ from the lattice Boltzmann and lattice gas systems to the theoretical $\nu_{\omega}$-value from Eq. (61) for a given value of $\omega$. Fig. 5 shows the results of this comparison for a high density ($\bar{N}^{eq}=1,000$) system.
Figure 6: The extracted amplitude of the decaying sound wave for both lattice gas and lattice Boltzmann simulations over 850 iterations run at $\omega=1$. along with two potential fittings at a constant decay rate of for the lattice gas. Simulations are initialized identically to those in Fig. 4, except with a $\bar{N}^{eq}$-value of 10. Both fits have been made utilizing non-linear curve fitting on xmgrace assuming Eq. (72) as the fitting equation, with the fits restricted to Iterations$<300$ (early-time) and $400<$Iterations$<700$ (late-time).
There is excellent agreement between the results of the lattice Boltzmann and lattice gas algorithms in this case, as well as between the results from both algorithms and Eq. (61). However, this is only true when considering high density systems. For low density systems, the decay rate is non-constant, an effect that can be seen in Fig. 6.
Both the lattice gas and its fittings show a clear distinction from the lattice Boltzmann in the case of this smaller $\bar{N}^{eq}$-value. The fits themselves are also imperfect, due to the fact that Eq. (72) assumes $\lambda$ to be a constant value. However because the early-time fitting and late-time fitting do not agree with each other, the value of $\lambda$ is not constant with respect to time for the lattice gas.
@@
We start by initializing the systems identically to those of Fig. 4, but with varying initial values of $\bar{N}^{eq}$. All systems were then run with $\omega=1$ for 1,000 timesteps, averaged over 100,000 seeds per density. The viscosities were calculated identically to those in Fig 5. Two sets of fitting parameters were used, corresponding to the early-time and late-time fittings in Fig. 6. We then repeated this process for the BGK entropic lattice Boltzmann algorithm.
We also considered the collision operator for an uncorrelated Poisson distribution. Since the slope of the collision operator from Fig. 2 (a) is slightly different to that of the BGK collision operator, we can obtain a different effective $\omega$-value with which to calculate the viscosity. However, as we will see below, this small correction does not explain the much larger deviations measured in our simulation.
@@ IV.2 The SOD Shock Tube
The SOD shock tube is defined by an initial state with two regions in equilibrium with differing initial average densities, but identical initial velocities Sod (1978). Due to the step discontinuity between the two regions, the SOD shock tube is a Riemann problem Toro (2009). This particular Riemann problem has been used as a basic test case for a variety of hydrodynamic simulation methods Sod (1978); Zha (1999); Loubère et al. (2014); Feng et al. (2010).
@@
$\bar{N}_{\mathrm{shock}}(x,0)=\begin{cases}N_{l}&x<\frac{L}{2}\\ N_{r}&\frac{L}{2}\leq x\\ \end{cases}$ (73)
with $N_{l}$ and $N_{r}$ being the average densities of the two initial regions. The initial $\tilde{f}_{i}$-values for the shock tube can then be written using Eq. (32), as both regions should initially be in equilibrium internally. The boundary conditions typically include reflective boundary conditions on both ends of the tube, however we found that a mirrored shock wave system allowed us to use periodic boundary conditions, though only one side of the shock wave will be shown due to their symmetry.
The analytical solution for the SOD shock tube for zero viscosity is relatively well studied as a solution to the Euler equations Toro (2009), however our collision operator restricts the system to an isothermal average state. Negro et al. derived an analytical solution for the isothermal inviscid SOD shock tube system Negro et al. (2019). For a rightwards moving shock ($N_{l}>N_{r}$) Negro’s analytical solutions for density and velocity can be written as,
$\bar{N}_{\mathrm{shock}}(x,t)=\begin{cases}N_{l}&x<x_{r}\\ N_{l}\exp\left(\frac{x_{0}-x}{tc_{s}}+1\right)&x_{r}\leq x<x_{c}\\ N_{l}\exp(-\zeta)&x_{c}\leq x<x_{s}\\ N_{r}&x\geq x_{s},\\ \end{cases}$ (74)
@@
$x_{s}=\frac{\zeta c_{s}t}{1-\frac{N_{r}}{N_{l}}\exp(\zeta)}+x_{0}.\$ (76)
@@
$02+\zeta^{2}-\frac{N_{l}}{N_{r}}\exp(-\zeta)-\frac{N_{r}}{N_{l}}\exp(\zeta)=$ 0 . (79)
@@
Figure 8: The $\bar{N}$ (a) and $\bar{J}$ (b) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=1$ comparing the lattice gas results averaged over 25,000 seeds and the lattice Boltzmann results. The system was initialized with a size of 150 lattice sites, and $\bar{N}=\rho$ values of $N_{l}=1,000$ and $N_{r}=500$, along with a mirrored shock wave on the opposite side of the periodic boundary. The lattice gas $n_{i}$-values are assigned utilizing a Poisson distribution around Eq. (32), with the $\bar{N}_{\mathrm{shock}}$-values specified previously and $u_{\mathrm{shock}}(x,0)=0$. (a) includes the analytical solution from Eq. (74). The analytical solution in (b) is the product of Eqs. (74) and (75). (a) also includes labels for various features of the shock wave.
@@
Fig. 8 shows the equivalence between the lattice Boltzmann and lattice gas simulations with the analytical solution after evolution. Specifically, it considers the ensemble averaged values of the $N$ and $J$ moments: $\bar{N}$ and $\bar{J}$ from Eqs. (49) and (50). $\bar{N}$ can be compared directly with $\bar{N}_{\mathrm{shock}}$ from the analytical solution in Eq. (74), and $\bar{J}$ can be compared to the analytical solutions in Eqs. (74) and (75) multiplied together, since $\bar{J}=u\bar{N}$. Fig. 8 (a) also labels the features of the SOD shock wave, including the rarefaction fan, the shock wave front, and the positions $x_{r}$, $x_{c}$, and $x_{s}$.
@@
The averaged $N$ (Fig. 8 (a)) and $J$ (Fig. 8 (b)) moments for the lattice gas algorithms show an excellent agreement with those of the lattice Boltzmann algorithm. They also show decent agreement with their respective analytical solutions for the inviscid SOD shock tube. There are some small deviations between both sets of simulated data and the analytical solution due to the viscosity not being low enough to recover the piecewise nature of the analytical solutions. For lattice Boltzmann simulations, it is possible to achieve closer agreement by using $\omega>1$, however, as mentioned above, collision probabilities greater than one are not yet an option for the hydrodynamic lattice gas algorithm.
@@
While gathering data for Fig. 8, we also considered the initial fluctuations in the moments, as we wished to see whether initial system fluctuations had any effect on the collision operator the way correlations were shown to in Fig. 2 (a). To this end, simulations with two different initial particle distributions were run. The first of these distributions is a Poisson distribution of the $n_{i}$-values given by Eq. (31), with Eq. (73) and $N_{l}=1,000,N_{r}=500$ for its initial $\bar{N}$-values. The second distribution was a flat distribution given by a delta function around the entropic lattice Boltzmann equilibrium distribution,
@@
Figure 9: Normalized pre- and post-collision $\pi$-values of the SOD Shock Tube. The systems were initialized identically to those in Fig. 8. After 75 iterations, another collision was run for each system and the pre- and post-collision $\pi$-values were measured for that extra collision. The lattice gas was averaged over 100,000 seeds. Both were normalized according to the average post-collision $\pi$-value from Eq. (70).
Since the system dynamics are driven by the collision operator we examined the pre- and post-collision $\bar{\pi}$-values for a single collision, similarly to what was done in Fig. 3. We can see in Fig. 9 that these pre- and post-collision results show good agreement between the lattice Boltzmann and lattice gas algorithms. We can also note the fact that for the parts of this system that are in equilibrium, the average collision operator is identical between the pre- and post-collision states. These equilibrium regions include both those that were initialized in equilibrium and have yet to be affected by the shock behavior, as well as the bulk of the shock wave itself ($x_{c}\leq x<x_{s}$). The other regions of the wave, including the rarefaction fan and the shock front, show highly non-equilibrium behavior. The shock front shows an increase in the average pre-collision $\pi$-value, and the rarefaction fan shows a decrease in the average pre-collision $\pi$-value.

### v2

@@ IV Non-equilibrium Dynamics
In this section we will apply our algorithm to two specific cases to see when the BGK approximation for the lattice Boltzmann limit of our lattice gas is an appropriate approximation, and in what cases we see divergent behavior. We will find that over-all the agreement is very good, but there clearly are some cases where significant deviations are observed. We then examine a strongly non-equilibrium shock wave and find surprisingly close agreement between the BGK LBM approximation and the Boltzmann limit of the lattice gas.
@@ IV.1 The Decaying Sound Wave
To test the approximate Boltzmann limit of our lattice gas, we start by analyzing the decaying isothermal sound wave. While the isothermal nature makes this a non-physical application, it is still valuable to compare the lattice Boltzmann and lattice gas algorithms without addressing the complication of including energy conservation. An analytical solution for this system was found by Luiz Czelusniak [33] in his thesis, and is given by,
@@
$\lambda=\left(\frac{2\tilde{\pi}}{L}\right)^{2}\nu$ (67)
@@
$f^{2}=\left(\frac{2\tilde{\pi}}{L}\right)^{2}\theta-\lambda^{2},$ (68)
@@
The sound wave must be initialized utilizing a particle distribution for lattice gas systems due to the discrete nature of particles. As the global equilibrium ensemble for the lattice gas is approximately Poisson distributed, we initialize our system by sampling the occupation numbers out of the Poisson distribution from Eq. (30), with $\tilde{f_{i}}$ values given by Eq. (31). The $\bar{N}$ and $u$ values needed to find the $\tilde{f_{i}}$ values can then be defined as the initial conditions for Eqs. (65) and (66) respectively,
@@
Figure 3: Normalized pre- and post-collision $\pi$ values of a decaying sound wave. The lattice Boltzmann system was initialized using Eq. (70) with $\bar{N}^{eq}=1,000$ and $L=250$, and the lattice gas used these same values as a mean for its Poisson distributed initialization. After 50 iterations run at $\omega=\omega_{LB}=1$
, the pre- and post-collision $\pi$ values were measured over the course of a single collision. The lattice gas was averaged over 100,000 random seeds. Both were normalized according to the equilibrium average post-collision $\pi$ value from Eq. (72).
We can observe the effect of a single collision on the system by looking at the difference between the pre- and post-collision averaged $\pi$ values of the system. We then normalize the system with the expected average post-collision $\pi$ value, which is found by using the $f_{i}^{0}$ from Eq. (31),
@@
To observe the viscosity of the system, we first consider the amplitude decay of the sine wave as it approaches equilibrium. We utilize a method of amplitude extraction discussed by Blommel and Wagner in an appendix to their paper [19]. Specifically, the amplitude of our sound wave is given by,
@@
Figure 4: The extracted amplitude evolution of the decaying sound wave for various relaxation times for the lattice gas and lattice Boltzmann simulations over 850 iterations for a lattice size of 50. Systems were initialized otherwise identically to those in Fig. 3, and run for $\omega=\omega_{LB}=1.0$, $\omega=\omega_{LB}=0.7$, and $\omega=\omega_{LB}=0.1$.
The amplitude was extracted utilizing Blommel and Wagner’s method after each iteration [19]. Lattice gas systems were averaged over 25,000 random seeds.
Fig. 4 shows the results of this amplitude extraction for $\omega=\omega_{LB}=1$, $\omega=\omega_{LB}=0.7$, and $\omega=\omega_{LB}=0.1$ over 850 iterations for an average of $\bar{N}^{eq}=1,000$ particles. There is excellent agreement between the lattice gas and the lattice Boltzmann algorithms in this case, despite the fact that the BGK lattice Boltzmann algorithm is not an exact Boltzmann limit for our lattice gas. We can measure the viscosity of the system numerically by extracting the decay rate from the wave amplitude. We can do this by fitting our results from Fig. 4 to the analytical solution from Eq. (65). The amplitude of Eq. (65) is given by,
@@
Figure 5: The recovered viscosities of the lattice Boltzmann and lattice Gas corresponding to various relaxation times along with the theoretical scaling from Eq. (63). Each system was initialized and run identically to those in Fig. 4. Along with the systems from Fig. 4, systems were also run at $\omega=\omega_{LB}=0.3$ and $\omega=\omega_{LB}=0.5$. The $\lambda$ values were found via non-linear curve fitting on xmgrace with the analytical solution from Eq. (74) and calculated using Eq. (67).
We can then compare the extracted values of $\nu$ from the lattice Boltzmann and lattice gas systems to the theoretical $\nu_{\omega}$ value from Eq. (63) for a given value of $\omega$. Fig. 5 shows the results of this comparison for a high density ($\bar{N}^{eq}=1,000$) system.
Figure 6: The extracted amplitude of the decaying sound wave for both lattice gas and lattice Boltzmann simulations over 850 iterations run at $\omega=\omega_{LB}=1$. along with two potential fittings at a constant decay rate of for the lattice gas. Simulations are initialized identically to those in Fig. 4, except with a $\bar{N}^{eq}$ value of 10. Both fits have been made utilizing non-linear curve fitting on xmgrace assuming Eq. (74) as the fitting equation, with the fits restricted to Iterations<300<300 (early-time) and 400<400<Iterations<700<700 (late-time).
There is excellent agreement between the results of the lattice Boltzmann and lattice gas algorithms in this case, as well as between the results from both algorithms and Eq. (63). The results from Figs. 4 and 5 imply that $\omega$ and $\omega_{LB}$ are equivalent, however this is only true when considering high density systems. For low density systems, the decay rate is non-constant, an effect that can be seen in Fig. 6.
Both the lattice gas and its fittings show a clear distinction from the lattice Boltzmann in the case of this smaller $\bar{N}^{eq}$ value. The fits themselves are also imperfect, due to the fact that Eq. (74) assumes $\lambda$ to be a constant value. However because the early-time fitting and late-time fitting do not agree with each other, the value of $\lambda$ is not constant with respect to time for the lattice gas.
@@
We start by initializing the systems identically to those of Fig. 4, but with varying initial values of $\bar{N}^{eq}$. All systems were then run with $\omega=1$ for 1,000 timesteps, averaged over 100,000 random seeds per density. The viscosities were calculated identically to those in Fig 5. Two sets of fitting parameters were used, corresponding to the early-time and late-time fittings in Fig. 6. We then repeated this process for the BGK entropic lattice Boltzmann algorithm for $\omega_{LB}=1$.
We also considered the collision operator for an uncorrelated Poisson distribution. Since the slope of the collision operator from Fig. 2 (a) is slightly different to that of the BGK collision operator, we can obtain a different effective $\omega$ value with which to calculate the viscosity. However, as we will see below, this small correction does not explain the much larger deviations measured in our simulation.
@@ IV.2 The SOD Shock Tube
The SOD shock tube is defined by an initial state with two regions in equilibrium with differing initial average densities, but identical initial velocities [34]. Due to the step discontinuity between the two regions, the SOD shock tube is a Riemann problem [35]. This particular Riemann problem has been used as a basic test case for a variety of hydrodynamic simulation methods [34, 36, 37, 38].
@@
[…]
with $N_{l}$ and $N_{r}$ being the average densities of the two initial regions. The initial $\tilde{f}_{i}$ values for the shock tube can then be written using Eq. (31), as both regions should initially be in equilibrium internally. The boundary conditions typically include reflective boundary conditions on both ends of the tube, however we found that a mirrored shock wave system allowed us to use periodic boundary conditions, though only one side of the shock wave will be shown due to their symmetry.
The analytical solution for the SOD shock tube for zero viscosity is relatively well studied as a solution to the Euler equations [35], however our collision operator restricts the system to an isothermal average state. Negro et al. derived an analytical solution for the isothermal inviscid SOD shock tube system [39]. For a rightwards moving shock ($N_{l}>N_{r}$) Negro’s analytical solutions for density and velocity can be written as,
[…]
@@
$x_{s}=\frac{\zeta c_{s}t}{1-\frac{N_{r}}{N_{l}}\exp(\zeta)}+x_{0}.\\$ (78)
@@
$2+\zeta^{2}-\frac{N_{l}}{N_{r}}\exp(-\zeta)-\frac{N_{r}}{N_{l}}\exp(\zeta)=0.$ (81)
@@
Figure 8: The $\bar{N}$ (a) and $\bar{J}$ (b) moment graphs of the isothermal SOD shock tube at 75 iterations for $\omega=\omega_{LB}=1$ comparing the lattice gas results averaged over 25,000 random seeds and the lattice Boltzmann results. The system was initialized with a size of 150 lattice sites, and $\bar{N}=\rho$ values of $N_{l}=1,000$ and $N_{r}=500$, along with a mirrored shock wave on the opposite side of the periodic boundary. The lattice gas $n_{i}$ values are assigned utilizing a Poisson distribution around Eq. (31), with the $\bar{N}_{\mathrm{shock}}$ values specified previously and $u_{\mathrm{shock}}(x,0)=0$. (a) includes the analytical solution from Eq. (76). The analytical solution in (b) is the product of Eqs. (76) and (77). (a) also includes labels for various features of the shock wave, and an inset that compares a single seed run of the lattice gas to the averaged lattice gas in the main body of (a). (b) also includes an inset that displays the Mach number, calculated using Eq. (82).
@@
Fig. 8 shows the equivalence between the lattice Boltzmann and lattice gas simulations with the analytical solution after evolution. Specifically, it considers the ensemble averaged values of the $N$ and $J$ moments: $\bar{N}$ and $\bar{J}$ from Eqs. (48) and (49). $\bar{N}$ can be compared directly with $\bar{N}_{\mathrm{shock}}$ from the analytical solution in Eq. (76), and $\bar{J}$ can be compared to the analytical solutions in Eqs. (76) and (77) multiplied together, since $\bar{J}=u\bar{N}$. Fig. 8 (a) also labels the features of the SOD shock wave, including the rarefaction fan, the shock wave front, and the positions $x_{r}$, $x_{c}$, and $x_{s}$, and includes an inset that shows the fluctuating nature of each run of the lattice gas. This inset shows the result of a run with no averaging, compared to our ensemble average lattice gas used in the main body of Fig 8 (a). Fig. 8 (b) includes an inset for the Mach number at each position along the shock wave. Here we calculate the Mach number assuming a constant speed of sound on our lattice given by Eq. (64),
$M=\frac{u}{c_{s}},$ (82)
however this only serves as an approximation, as $c_{s}$ is potentially velocity dependent, as was discussed briefly earlier.
@@
The averaged $N$ (Fig. 8 (a)) and $J$ (Fig. 8 (b)) moments for the lattice gas algorithms show an excellent agreement with those of the lattice Boltzmann algorithm. They also show decent agreement with their respective analytical solutions for the inviscid SOD shock tube. There are some small deviations between both sets of simulated data and the analytical solution due to the viscosity not being low enough to recover the piecewise nature of the analytical solutions. For lattice Boltzmann simulations, it is possible to achieve closer agreement by using $\omega_{LB}>1$, however, as mentioned above, collision probabilities greater than one are not yet an option for the hydrodynamic lattice gas algorithm.
@@
While gathering data for Fig. 8, we also considered the initial fluctuations in the moments, as we wished to see whether initial system fluctuations had any effect on the collision operator the way correlations were shown to in Fig. 2 (a). To this end, simulations with two different initial particle distributions were run. The first of these distributions is a Poisson distribution of the $n_{i}$ values given by Eq. (30), with Eq. (75) and $N_{l}=1,000,N_{r}=500$ for its initial $\bar{N}$ values. The second distribution was a flat distribution given by a delta function around the entropic lattice Boltzmann equilibrium distribution,
@@
Figure 9: Normalized pre- and post-collision $\pi$ values of the SOD Shock Tube. The systems were initialized identically to those in Fig. 8. After 75 iterations, another collision was run for each system and the pre- and post-collision $\pi$ values were measured for that extra collision. The lattice gas was averaged over 100,000 random seeds. Both were normalized according to the average post-collision $\pi$ value from Eq. (72).
Since the system dynamics are driven by the collision operator we examined the pre- and post-collision $\bar{\pi}$ values for a single collision, similarly to what was done in Fig. 3. We can see in Fig. 9 that these pre- and post-collision results show good agreement between the lattice Boltzmann and lattice gas algorithms. We can also note the fact that for the parts of this system that are in equilibrium, the average collision operator is identical between the pre- and post-collision states. These equilibrium regions include both those that were initialized in equilibrium and have yet to be affected by the shock behavior, as well as the bulk of the shock wave itself ($x_{c}\leq x<x_{s}$). The other regions of the wave, including the rarefaction fan and the shock front, show highly non-equilibrium behavior. The shock front shows an increase in the average pre-collision $\pi$ value, and the rarefaction fan shows a decrease in the average pre-collision $\pi$ value.